Listed below are the overhead widths​ (in cm) of seals measured from photographs and the weights​ (in kg) of the seals. Construct a​ scatterplot, find the value of the linear correlation coefficient​ r, and find the critical values of r using

alphaαequals=0.01

Is there sufficient evidence to conclude that there is a linear correlation between overhead widths of seals from photographs and the weights of the​ seals?

Overhead Width: 7.2, 7.6, 9.7, 9.2, 8.8, 8.2

Weight: 117, 197, 242, 197, 204, 287

Construct a scatter plot of the data:

What is the linear correlation coefficient r?:

What is the critical value:

Is the absolute value of the linear correlation coefficient less or greater than the positive critical value, and is there sufficient evidence to support the claim that there is a linear correlation between overhead widths from photographs and the weights of the seals for a significance level of 0.01?
Answer #1

Solution:

Find X.Y,X2 and Y2 as it was done in the table below 51.84 13689 7.2 842.4 7.6 197 1497.2 57.76 38809 9.7 242 2347.4 94.09 58

df = n - 2 = 6-2 = 4, look for critical values in r- table, 0.917 for α = 0.05. We take both signs for 2-tailed test
The critical values are r = -0.917, 0.917

Because the absolute value of the linear correlation coefficient is less than the positive critical value, we fail to reject null hypothesis, i.e., and there is not sufficient evidence to support the claim that there is a linear correlation between overhead widths from photographs and the weights of the seals for a significance level of 0.01